# Smooth-edged Perturbations Improve Perturbation-based Image Explanations

### v1

@@ Abstract
Perturbation-based post-hoc image explanation methods are commonly used to explain image prediction models by perturbing parts of the input to measure how those parts affect the output. Due to the intractability of perturbing each pixel individually, images are typically attributed to larger segments. The Randomized Input Sampling for Explanations (RISE) method solved this issue by using smooth perturbation masks.
While this method has proven effective and popular, it has not been investigated which parts of the method are responsible for its success. This work tests many combinations of mask sampling, segmentation techniques, smoothing, and attribution calculation. The results show that the RISE-style pixel attribution is beneficial to all evaluated methods. Furthermore, it is shown that attribution calculation is the least impactful parameter. The implementation of this work is available online  [^1] https://github.com/guspih/post-hoc-image-perturbation.
@@ I Introduction
Some research in XAI focuses on developing models that are inherently explainable [3]. Other research uses so-called global methods that attempt to explain the entirety of a model’s prediction space [4]. However, these approaches are not suitable for DNNs. A popular approach that avoids these problems is post-hoc explanations [5].
Post-hoc explanations forego trying to understand the model in its entirety and focus instead on explaining individual predictions. For example, instead of explaining the entire process by which a bank makes loan decisions the banker only needs to explain the parts of the process that are important for a given decision. One type of post-hoc explanation that is popular in the computer vision domain is perturbation-based explanations. Perturbation-based explanations work by analyzing how the model’s predictions change as the original input is perturbed. As they only need the given inputs and outputs perturbation-based explanations are model-agnostic and can be applied to any model.
@@
Figure 1: The pipeline for perturbation-based image attribution used in this work. The image is segmented, samples indicating what segments to perturb are drawn, the sampled segments are perturbed, the model to explain makes predictions for the perturbed samples, and the input-output pairs are used to calculate attribution per-segment and per-pixel.
@@
Since attribution is calculated based on which segments are perturbed, most methods assign attribution per-segment, but cannot differentiate the influence between pixels. The Randomized Input Sampling for Explanations (RISE) method solves this by using smooth perturbations, where pixels are perturbed more as they get closer to the segment center [9]. This is then used to calculate a per-pixel attribution by weighing the attribution of a pixel by how perturbed the pixel was.
Like many perturbation-based explanations, RISE is introduced as an entire pipeline from segmenting to attribution. This work explores how the benefits of smooth-edged perturbations can benefit other perturbation-based pipelines. It also expands on the original RISE implementation by evaluating a variety of segmentation, sampling, perturbation, and attribution methods using occlusion metrics [8]. The evaluations are carried out on the ImageNet validation set [13] for three different CNNs [14, 15, 16] using both per-segment and per-pixel attributions. The different pipeline parameters that have been combined and evaluated are shown in Table I.
The results show that using smooth edges and weighing pixels by how faded they are in a given sample improves the performance of all evaluated methods. Another noteworthy result is that the method of calculating the attribution, which is typically what is highlighted as the most important part, has little impact on performance. Conversely, the sampling, number of samples, segmentation, and per-pixel attribution all have a greater impact on performance.
@@
Grid+Bilinear Random 4000/8000 AlexNet CIU
Grid+Gaussian Entropic 400 VGG-16 PDA
SLIC+Gaussian Only one 50 ResNet LIME
All but one SHAP
@@ II Methodology
This work evaluates pipelines using all possible combinations of the different segmenting, sampling, perturbing, and attribution methods as well as sample sizes listed in Table I. Each pipeline is tested with three different ImageNet [13] pretrained CNNs by using them to explain the models’ predictions on the ImageNet validation set and then evaluating those explanations using occlusion metrics [8]. The different parts of the experiments are described in detail in the following subsections.
@@ II-A Segmenting
The experiments use the same $7\times 7$ grid of segmentation as the original RISE implementation [9]. To make the SLIC segmentation as similar to the grid implementation as possible, SLIC is instantiated with 49 segment centers in the experiments. The default scikit-image implementation for SLIC is used [18].
@@ II-B Sampling
This work generates samples indicating which segments to perturb using random, entropic, "only one", and "all but one" sampling. Random sampling consists of, for each segment and sample, randomly deciding whether it should be perturbed with a probability $p$. In this work $p=0.5$.
Entropic sampling is created to be similar to the default KernelSHAP sampling behavior [12]. Entropic sampling will first sample the low-entropy samples, i.e., samples with as many or as few segments perturbed as possible. No segments are perturbed in the first sample, all segments are perturbed in the second, followed by all possible combinations of one segment perturbing and all combinations of one segment unperturbed, followed by combinations of two segments perturbed/unperturbed, and so on.
"Only one" and "All but one" sampling consists of creating all samples where only one segment is perturbed and where all but one segment is perturbed respectively. Both methods also add the sample where no segments are perturbed as this is needed by the Contextual Importance and Utility (CIU) attribution [19].
Random and entropic sampling are evaluated for three different sample sizes.The 4000/8000 sample size is used to be consistent with the original RISE evaluation. AlexNet and VGG-16 use 4000 samples and ResNet models were evaluated with 8000 samples. A sample size of 50 is used with all four methods, where "only one" and "all but one" sampling will always create one more sample than the number of segments.
@@ II-C Perturbing
Perturbing consists of pixel-wise multiplication between the normalized image and a perturbation mask of values between 0 and 1. The mask is created by setting all values in the segments to be perturbed to 0 and all others to 1, then the mask is smoothened so that the values closer to the center of each segment are close to 0 and those at the edges and beyond are closer to 1. Thus pixels outside the perturbed segments are mostly unchanged, but fade towards the normalization mean as they get closer to the segment centers. The original implementation achieves this by using bilinear upsampling to scale a $7\times 7$ grid of 0s and 1s to the size of the full image, an implementation that is replicated in this work. An issue with this method is that it relies on having a lower resolution mask to upscale which excludes using some popular segmentation methods such as SLIC. To combat this issue another method of creating smooth segment masks through applying a Gaussian filter is introduced. For this work, the Gaussian filter has a $\sigma=10$ which gives similar masks when compared to bilinear upscaling.
@@ II-D Attributing
This work evaluates five existing attribution methods, CIU [19], PDA [10], LIME [11], SHAP [12], and RISE [9]. Some of these methods cover more parts of the pipeline than just attribution. However, in this work, the method names are used as a shorthand for the attribution calculation from the input-output pairs created by the predicting step of the pipeline.
@@
Prediction Difference Analysis (PDA) [10] works similarly to CIU, but uses average difference instead of maximum difference. PDA has been adapted to work with images [10], though both in the original and image implementation only a single feature is changed at a time. In this work, PDA has been generalized to work when multiple features are perturbed simultaneously. The PDA attribution is given by $w_{PDA}(s)=Y-avg(Y\backslash_{s})$.
@@
The attribution used by RISE [9] is similar to PDA, but instead of using the average decrease when the feature is perturbed, it uses the average prediction when it is not perturbed. RISE attribution for a segment is given by $w_{RISE}(s)=avg(Y\backslash_{\bar{s})}$.
@@
Additionally, RISE attribution utilizes smooth pixel perturbation masks to assign per-pixel attributions according to $w_{p}=\frac{1}{\sum_{s\in S}M_{s}^{p}}\sum_{s\in S}w_{s}\cdot M_{s}^{p}$, where $M_{s}^{p}$ is the value of pixel $p$ in the perturbation mask of segment $s$. Note that this calculation means that pixels outside segment $s$ which were slightly perturbed due to the smooth mask, also include that influence in the calculation. For example, this means that pixels at segment borders get a lesser influence from both segments. This work evaluates attribution both per-segment ($w_{s}$) and per-pixel ($w_{p}$).
II-E Evaluation
@@
The various pipelines are tested by explaining the predictions of three ImageNet pretrained CNNs on the ImageNet validation set and evaluating those explanations with occlusion metrics. The three pretrained CNNs are AlexNet [14, 23], VGG-16 [15], and ResNet-50 [16] using trained parameters from the Torchvisio $0.15.2$ framework [24]. The input to the models is normalized using the average pixel values of ImageNet.
@@
Evaluation is carried out using one image per class of the ImageNet validation for a total of 1000 images ($2\%$ of the total validation set). Limited evaluation was performed on the full validation set and no statistically significant ($p<0.05$) difference could be found compared to using 2% of the data. For each image, the top predicted class of each model was explained through segment and pixel attribution using each pipeline. The attribution was then evaluated using occlusion metrics. The occlusion metrics used in this work are similar to the ones used for evaluating the original RISE implementation [9] though modified to take advantage of recent findings that increase the consistency [8].
@@
Occlusion metrics consist of increasingly occluding the image and observing how the prediction changes. By occluding the pixels with the Least Influence First (LIF), the model prediction is expected to be similar until the influential pixels start getting occluded. Conversely, by occluding the pixels with the Most Influence First (MIF), the model prediction is expected to lower quickly. A good explanation should have a large area under the LIF prediction-occlusion curve and a small area under the MIF curve. LIF and MIF are equivalent to the insertion and deletion metrics used to evaluate RISE originally. The LIF and MIF metrics are highly variable but the combined metric ($LIF-MIF$), called Symmetric Relevance Gain (SRG), is more reliable [8]. The connection between the three metrics is visualized in Fig. 2.
@@
This work uses the SRG metric to evaluate performance. It is calculated by occluding the image over a total of 10 equal steps (from 0% occlusion in step 1 to 100% occlusion in step 10). The remaining pixels with the lowest or greatest attribution score for original top-class prediction are occluded in each step for LIF and MIF respectively. When there are many pixels with the same attribution, then pixels are chosen in an arbitrary deterministic order. Occlusion is performed by setting the pixels to the mean pixel value of the image, which mirrors one of the evaluation methods explored by Blücher et al., [8]. The average of the original top-class prediction over these 10 images is then recorded as the LIF and MIF scores. The SRG score is calculated as $LIF-MIF$.
@@ III Results and Analysis
The results of different combinations of segmenting, perturbing, and attribution as the average SRG metric can be found in Table II. Notably, for all combinations of segmenting, perturbing, and attribution methods using per-pixel instead of per-segment attribution improves performance. Furthermore, the improvement of using per-pixel rather than per-segment is significantly greater than switching attribution methods. Using a Gaussian filter instead of bilinear upsampling does not affect performance, except for a mild increase in SRG. SLIC performs much better than Grid segmenting in all cases but sees a relatively smaller improvement when using per-pixel attribution. This is likely due to SLIC having better boundaries between segments.
TABLE II: The average SRG in % for all pipelines with different combinations of segmenting, perturbing, and attribution methods with either per-segment or per-pixel attribution.
@@
The average SRG for pipelines with different sampling methods and sample sizes over the different attribution methods is shown in Table III. Unsurprisingly, increasing sample size yields improved performance. What is surprising is that random sampling significantly outperforms entropic and does so even for SHAP for which it is specifically adapted. PDA struggles with entropic sampling, except for when the sample size is 50, which is almost equivalent to "only one" sampling. Again it is noteworthy that the attribution method is the least impactful factor, except under some combinations of sampling and sample size that seem to cause some attribution methods to fail. All attribution have the same performance for Entropic, "only one" and "all but one" sampling with a sample size of 50 as under these limited circumstances the order of influential segments is equivalent for CIU, PDA, and RISE. At the same time, LIME and SHAP converge to the same ordering.
TABLE III: The average SRG in % for all pipelines with different combinations of sampling and attribution methods.
Sampling Sample size CIU PDA LIME SHAP RISE
Random 4000/8000 - 25.6 25.8 24.0 25.6
Entropic 4000/8000 - 14.7 18.2 18.8 17.8
[…]
Entropic 400 - 9.0 15.6 17.3 15.0
[…]
Entropic 50 - 13.3 13.3 13.3 13.3
Only one 50 13.3 13.3 13.3 13.3 13.3
All but one 50 14.9 14.9 14.9 14.9 14.9

### v2

# Segmentation and Smoothing Affect Explanation Quality More Than the Choice of Perturbation-based XAI Method for Image Explanations

@@ Abstract
Perturbation-based post-hoc image explanation methods are commonly used to explain image prediction models. These methods perturb parts of the input to measure how those parts affect the output. Since the methods only require the input and output they can be applied to any model, making them a popular choice to explain black-box models. While many different models exist and have been compared with one another, it remains poorly understood which parameters of the different methods are responsible for their varying performance.
This work uses the Randomized Input Sampling for Explanations (RISE) method as a baseline to evaluate many combinations of mask sampling, segmentation techniques, smoothing, attribution calculation, and per-segment or per-pixel attribution, using a proxy metric. The results show that attribution calculation, which is frequently the focus of other works, has little impact on the results. Conversely, segmentation and per-pixel attribution, rarely examined parameters, have a significant impact.
The implementation of and data gathered in this work are available online  [^1] https://github.com/guspih/post-hoc-image-perturbation [^2] https://bit.ly/smooth-mask-perturbation.
@@ I Introduction
The complexity of DNNs makes it difficult to explain the model in its entirety. Instead, methods for explaining DNNs tend to focus on explaining individual predictions made by the model [3]. Some research focuses on developing inherently interpretable models [4]. However, these methods are bound to specific models and cannot be applied to the vast amount of non-interpretable models that exist and are being developed [3]. In these cases, it is common to use post-hoc explanations that are separate from the models and can be applied during prediction [5].
Many different types of post-hoc methods exist to explain computer vision models. Though many methods can only be used on models that satisfy certain criteria, such as differentiability in the case of gradient-based methods. This work instead concerns a type of post-hoc explanation method that does not place any restriction on the models being used; so-called perturbation-based methods. Perturbation-based explanations work by analyzing how the model’s predictions change as the original input is perturbed. For image prediction, this is commonly used to find the regions of the image that cause the model to make a certain prediction. This can be used to assess if a specific model or prediction is sound and help locate important features for a user.
@@
Figure 1: The pipeline for perturbation-based image attribution used in this work. The image is segmented, samples indicating what segments to perturb are drawn, the sampled segments are perturbed, the model to explain makes predictions for the perturbed samples, and the input-output pairs are used to calculate per-segment and per-pixel attribution.
@@
The influence attributed to each segment is often visualized using attribution maps. Attribution maps are heat maps showing high influence using one color and fading to another color to show low influence. By analyzing the attribution map a human can understand the model’s behavior. For example, a prediction with a high influence on a non-salient part of the image can indicate problems with the model. Using attribution map explanations increases the joint human-AI performance on some tasks [13].
Since influence is attributed based on which segments are perturbed, most perturbation-based methods assign attribution to entire segments, but cannot differentiate the influence between pixels within a segment. The Randomized Input Sampling for Explanations (RISE) method solves this by using smooth perturbations, where pixels closer to the segment center are perturbed more [9]. With this approach, a per-pixel attribution can be calculated by assigning each pixel the influence of each segment weighted by how perturbed the pixel was for that segment.
As for most papers on perturbation-based explanations for images, in [9] RISE is described as an entire pipeline from segmenting to pixel attribution. As such it is not clear which parts of the RISE method, such as per-pixel attribution, that is responsible for its performance. This work explores the individual parameters of the perturbation-based pipeline to evaluate which parameters have the most impact on performance. This is done by taking a version of the original RISE pipeline and evaluating the original parameters alongside a variety of additional segmentation, sampling, perturbation, and attribution methods.
Due to the large number of parameter combinations to evaluate, human evaluation is not feasible for this study. Instead, this work uses a faithfulness metric dubbed Symmetric Relevance Gain (SRG) [8]. The evaluations are carried out on the ImageNet validation set [14] for three different CNNs [15, 16, 17] using both per-segment and per-pixel attributions. The different pipeline parameters that have been combined and evaluated are shown in Table I.
The results show that the method of calculating the attribution, which is typically what is highlighted as the most important part, has little impact on performance. Instead, segmentation and sampling, which are often overlooked, have a significant impact on performance. This calls into question studies that compare the original implementations of attribution calculation methods without correcting for the other parameters of those implementations. Additionally, per-pixel attribution gives a significantly higher average performance compared to per-segment attribution.
I-A Contributions
The primary contributions of this work are as follows.
  • Evaluating the different parameters of the perturbation-based image attribution pipeline finding that all parameters can have a significant impact on performance.
  • Adapting the smooth-mask enabled per-pixel attribution used in RISE to work with other perturbation-based attribution methods.
  • Improving the RISE pipeline by using Gaussian filtering instead of bilinear upsampling, which enables the use of many other segmenting methods.
  • Generalizing the Probabilistic Difference Analysis (PDA) [10] method to work with multiple perturbing multiple segments per sample.
@@ I-A Contributions
Grid+Bilinear Only one 4000/8000 AlexNet CIU
Grid+Gaussian All but one 400 VGG-16 PDA
SLIC+Gaussian Random 50 ResNet LIME
Entropic SHAP
@@ II Methodology
This work evaluates different segmenting, sampling, perturbing, and attribution methods using three different sample sizes as listed in Table I. All possible pipelines combining one of each parameter are evaluated by explaining the predictions made by three different ImageNet [14] pretrained CNNs as measured by the SRG metric [8]. The different parts of the experiments are described in detail in the following subsections.
@@ II-A Segmenting
The experiments use the same $7\times 7$ segmentation grid as the original RISE implementation [9]. To make the SLIC segmentation as similar to the grid implementation as possible, SLIC is instantiated with 49 segment centers in the experiments. The default scikit-image implementation for SLIC is used [19].
@@ II-B Sampling
This work generates samples indicating which segments to perturb using “only one”, “all but one”, random, and entropic sampling. “Only one” sampling generates all samples where only one segment is perturbed. “All but one” sampling generates all samples where all but one segment is perturbed. Both “Only one” and “All but one” sampling also generate the sample where no segment is perturbed as a reference value. Random sampling generates a given number of samples where each segment is perturbed with a probability of $0.5$. Entropic sampling generates a given number of samples in increasing order of entropy, i.e., samples with as many or as few segments perturbed as possible.
Entropic sampling is implemented to be similar to the default KernelSHAP sampling behavior [12]. No segments are perturbed in the first sample, all segments are perturbed in the second, followed by all possible combinations of one segment perturbing and all combinations of one segment unperturbed, followed by combinations of two segments perturbed/unperturbed, and so on until the sample size is reached.
“Only one” and “all but one” sampling create one sample per segment, which makes them computationally efficient. They are presumably the simplest sampling methods that can be used and have worked well with Contextual Importance and Utility (CIU) method for detecting bleeding in gastro-enterological [20]. Random and entropic sampling can generate samples consisting of combinations of segments, which can yield many more unique samples. Using combinations of segments presumably allows for taking into account inter-dependencies between different segments, potentially leading to more informative explanations.
II-B1 Sample size
Random and entropic sampling use sample sizes of 4000/8000, 400, as well as 50. The 4000/8000 sample size is used to be consistent with the original RISE evaluation. AlexNet and VGG-16 use 4000 samples and ResNet models were evaluated with 8000 samples. “Only one” and “all but one” sampling always generate samples equal to the number of segments plus one which is 50 samples at most in this work.
@@ II-C Perturbing
This work perturbs images by replacing the perturbed segments with a solid color, which is the most common way of perturbing images. The color used for perturbing is the mean color of the ImageNet training set, which is the color that will be normalized to 0 in the preprocessing of the Torchvision models used in this work, which is the same as the original RISE implementation. Like the original RISE implementation, all pixels in the segments are not replaced with exactly the given color and instead are faded between their original color and the normalization mean according to a segment mask that fades closer to the segment border. This mask is also later reused to attribute per-pixel in addition to the normal per-segment attribution.
Perturbing consists of pixel-wise multiplication between the normalized image and a perturbation mask of values between 0 and 1. The mask is created by setting all values in the segments to be perturbed to 0 and all others to 1, then the mask is smoothened so that the values closer to the center of each segment are close to 0 and those at the edges and beyond are closer to 1. Thus pixels outside the perturbed segments are mostly unchanged, but fade towards the normalization mean as they get closer to the segment centers. The original implementation achieves this by using bilinear upsampling to scale a $7\times 7$ grid of 0s and 1s to the size of the full image, an implementation that is replicated in this work. An issue with this method is that it relies on having a lower resolution mask to upscale which excludes using some popular segmentation methods such as SLIC. To combat this issue, this work introduces the use of Gaussian filters for mask smoothing. The Gaussian filter smoothens the mask by assigning each pixel in the mask the average value of all other pixels, weighted by a 2D Gaussian function centered on the pixel. This work uses a Gaussian filter with $\sigma=10$ which in this use case gives similar smooth masks when compared to bilinear upscaling.
@@ II-D Attributing
This work evaluates five existing attribution methods, CIU [21], PDA [10], LIME [11], SHAP [12], and RISE [9]. In this work, the method names refer only to the actual attribution calculation of each method.
@@
The Prediction Difference Analysis (PDA) method [10] works similarly to CIU, but uses average difference instead of maximum difference. PDA has been adapted to work with images [24], though both in the original and image implementation only a single feature is changed at a time (“only one” sampling). In this work, PDA has been generalized to work when multiple features are perturbed simultaneously. The PDA attribution is given by $w_{PDA}(s)=Y-avg(Y\backslash_{s})$.
@@
The attribution used by RISE [9] is similar to PDA, but instead of using the average decrease when the feature is perturbed, it uses the average prediction when it is not perturbed. This similarity means that in the case where every possible sample is used, the two methods have equivalent ranks of the segments by influence, which in turn produces equivalent SRG scores. RISE attribution for a segment is given by $w_{RISE}(s)=avg(Y\backslash_{\bar{s}})$.
@@
In addition to per-segment ($w_{s}$) attribution, this work also evaluates attribution per-pixel similar to the original RISE implementation. Typically, perturbation-based explanations assign the same attribution to all pixels in the same segment ($w_{s}$). However, by using the smoothing masks ($M_{s}$) used for perturbation the attribution scores of a pixel can instead be set to the average of the attribution of each segment weighted by how perturbed the pixel was by that segment. This is formalized as $w_{p}=\frac{1}{\sum_{s\in S}M_{s}^{p}}\sum_{s\in S}w_{s}\cdot M_{s}^{p}$, where $M_{s}^{p}$ is the value of pixel $p$ in the perturbation mask of segment $s$. Note that this calculation means that pixels outside segment $s$ which were slightly perturbed due to the smooth mask, also include that influence in the calculation. This results in pixels at segment centers getting almost all their attribution score from that segment, while pixels at the borders of segments get a lesser attribution from all the segments they border.
II-E Evaluating
@@
The various pipelines are tested by explaining the predictions of three ImageNet pretrained CNNs on the ImageNet validation set [14] and evaluating those explanations using the Symmetric Relevance Gain (SRG) metric [8]. The three pretrained CNNs are AlexNet [15, 26], VGG-16 [16], and ResNet-50 [17] using trained parameters from the Torchvision $0.15.2$ framework [27]. The input to the models is normalized using the average pixel values of ImageNet.
@@
The ImageNet validation set is used as it is the same dataset on which RISE was originally evaluated. Evaluation is carried out using the first image of each class of the ImageNet validation for a total of 1000 images ($2\%$ of the total validation set). The limitation to 1000 images is to make the computation demands feasible as over 300 pipelines were evaluated, some requiring 8000 model calls per image. Despite the limited number of images and some pipelines using the same model calls as they share the first parameters, the evaluation still required close to a billion model calls. Limited evaluation was performed on the full validation set and no statistically significant ($p<0.05$) difference could be found compared to using 2% of the data. For each image, the top predicted class of each model was explained through segment and pixel attribution using each pipeline.
@@
The SRG metric is used due to the large number of parameter combinations to evaluate, which signifies that human evaluation is not feasible to use for this study. SRG is measured by increasingly occluding the image and observing how the prediction changes. By occluding the pixels with the Least Influence First (LIF), the model prediction is expected to be similar until the influential pixels start getting occluded. Conversely, by occluding the pixels with the Most Influence First (MIF), the model prediction is expected to lower quickly. A good explanation should have a large area under the LIF prediction-occlusion curve and a small area under the MIF curve. LIF and MIF are equivalent to the insertion and deletion metrics used to evaluate RISE originally [9]. However, LIF and MIF have been shown to be inconsistent at ranking explanations as the occlusion type changes (e.g. random color, mean color, inpainting). The SRG metric, calculated as the difference between LIF and MIF, is more reliable [8]. This also eliminates some of the effects of occluded images being out-of-distribution. The connection between LIF, MIF, and SRG is visualized in Fig. 2.
@@
In this work, SRG is calculated by occluding the image over a total of 10 equal steps (from 0% occlusion in step 1 to 100% occlusion in step 10). The remaining pixels with the lowest or greatest attribution score for original top-class prediction are occluded in each step for LIF and MIF respectively. When there are many pixels with the same attribution, then pixels are chosen in an arbitrary deterministic order. Occlusion is performed by setting the pixels to the mean pixel value of the image, which mirrors one of the methods used for evaluating RISE originally [9]. The average of the original top-class prediction over these 10 images is then recorded as the LIF and MIF scores. The SRG score is calculated as $LIF-MIF$.
@@ III Results and Analysis
The results consist of the LIF, MIF, and SRG metrics for every attribution pipeline. As this is too much data to present in this work, it is summarized as the average SRG metric for different parameter combinations. The tables show for some parameter combinations the SRG in % averaged over all pipelines with those combinations (i.e., the average result for all combinations of the unspecified parameters). The complete data is available online [^3] https://bit.ly/smooth-mask-perturbation.
The average SRG for pipelines with different sampling methods and sample sizes over the different attribution methods is shown in Table II. Unsurprisingly, increasing sample size yields improved performance. What is surprising is that random sampling significantly outperforms entropic and does so even for SHAP, for which entropic is the default sampling method. It is noteworthy that the attribution method (row-wise) in general has less impact on the SRG value than the sampling method (column-wise). Furthermore, some attribution methods seem to fail when used together with certain sampling methods and sample sizes. For non-random sampling with a sample size of 50 all attribution methods have the same performance. Under these circumstances, CIU, PDA, and RISE mathematically produce the same rankings of segments to which LIME and SHAP consistently converge.
TABLE II: The SRG in % for all pipelines with different combinations of sampling and attribution methods. The results are averaged for all pipelines with the specified parameters.
Attribution
Sampling Sample size CIU PDA LIME SHAP RISE
Only one $\leq 50$ 13.3 13.3 13.3 13.3 13.3
All but one $\leq 50$ 14.9 14.9 14.9 14.9 14.9
Random 50 - 16.0 6.8 6.8 16.0
Entropic 50 - 13.3 13.3 13.3 13.3
[…]
Entropic 400 - 9.0 15.6 17.3 15.0
[…]
Entropic 4000/8000 - 14.7 18.2 18.8 17.8
The results of different combinations of segmenting, perturbing, and attribution can be found in Table III. Again, the attribution calculation method has less impact on performance compared to the other parameters. Notably, average performance increases when attributing per-pixel for all combinations of segmenting, perturbing, and attribution methods. The Gaussian filter does not give a significantly different performance for grid segmentation compared to bilinear upsampling. However, as the Gaussian filter was implemented to mimic bilinear upsampling in this case, the results show that there is no harm in choosing one over the other. Instead, the benefit of smoothing with a Gaussian filter becomes obvious in the significant performance increase seen with SLIC over grid segmentation. As SLIC with pixel attribution is not possible with bilinear upsampling, this suggests that Gaussian filters enable the use of better segmentation methods without sacrificing performance on grid segmentation.
TABLE III: The SRG in % for all pipelines with different combinations of segmenting, perturbing, and attribution methods with either per-segment or per-pixel attribution. The results are averaged for all pipelines with the specified parameters.
Attribution